Samples of 4 cards are drawn from a population of 6 cards numbered 1-6. Construct a sampling distribution of the sample means and answer the following questions:

1. How many samples of size 4 can be drawn from the population?

2. What are the possible means?

3. What is the probability of getting 4 as a mean?

4. What is the probability of getting 3.5 as a mean?


1
Expert's answer
2021-06-03T12:52:38-0400

We have population values "1,2,3,4,5,6" population size "N=6" and sample size "n=4." Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{6}{4}=15""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 1,2,3,4 & 2.5 \\
 \hdashline
 2 & 1,2,3,5 & 2.75 \\
 \hdashline
 3 & 1,2,3,6 & 3 \\
 \hdashline
 4 & 1,2,4,5 & 3 \\
 \hdashline
 5 & 1,2,4,6 & 3.25 \\
 \hline
 6 & 1,2,5,6 & 3.5 \\
 \hline
 7 & 1,3,4,5 & 3.25 \\
 \hline
 8 & 1,3,4,6 & 3.5 \\
 \hline
 9 & 1,3,5,6 & 3.75 \\
 \hline
 10 & 1,4,5,6 & 4 \\
 \hline
 11 & 2,3,4,5 & 3.5 \\
 \hline
 12 & 2,3,4,6 & 3.75 \\
 \hline
 13 & 2,3,5,6 & 4 \\
 \hline
 14 & 2,4,5,6 & 4.25 \\
 \hline
 15 & 3,4,5,6 & 4.5 \\
 \hline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 10/4 & 1& 1/15 & 9/20 & 81/80 \\
 \hdashline
 11/4 & 1 & 1/15 & 11/60 & 121/240 \\
 \hdashline
 12/4 & 2 & 2/15 & 24/60 & 288/240 \\
 \hdashline
 13/4 & 2 & 2/15 & 26/60 & 338/240 \\
 \hdashline
 14/4 & 3& 3/15 & 42/60 & 588/240 \\
 \hdashline
 15/4 & 2 & 2/15 & 30/60 & 450/240 \\
 \hdashline
 16/4 & 2 & 2/15 & 32/60 & 512/240 \\
 \hdashline
 17/4 & 1 & 1/15 & 17/60 & 289/240 \\
 \hdashline
 18/4 & 1 & 1/15 & 18/60 & 324/240 \\
 \hdashline
 Total & 15 & 1 & 207/60 & 2910/240 \\ \hline
\end{array}"



2.

Possible means "2.5,2.75,3,3.5,3.75,4,4.25,4.5"


3.


"P(\bar{X}=4)=\dfrac{2}{15}"

4.


"P(\bar{X}=3.5)=\dfrac{3}{15}"